OpenStudy (anonymous):

Suppose that you took out a loan with a principal of $24,680. To pay off the principal and the interest, you made quarterly payments of $1,382 for six years. Furthermore, you paid a service charge of $396. What was your total finance charge?

OpenStudy (phi):

If you make quarterly payments for 6 years, how many payments is that ?

OpenStudy (anonymous):

thats every 3 ,months right?

OpenStudy (phi):

yes

OpenStudy (anonymous):

288

OpenStudy (phi):

quarterly means 4 times a year (every 3 months, as you said) if you make 4 payments in one year, how many in 6 years ?

OpenStudy (anonymous):

ohh lol 24

OpenStudy (phi):

yes 24 payments (I am wondering what logic got you to 288... I hope you fix the "bug" in your thinking on that) if you make 24 payments, and each payment is $1,382 how much money do you pay out all together ?

OpenStudy (anonymous):

$33168

OpenStudy (anonymous):

plus service charge = 33564

OpenStudy (anonymous):

haha lol i did six times four times twelve lol

OpenStudy (anonymous):

yeah 33168 plusservice charge

OpenStudy (phi):

ok. total out is 33564 the finance charge is the amount over the original $24,680

OpenStudy (anonymous):

so i add those numbers ?

OpenStudy (phi):

think about it. If you borrowed 24680 and then paid back 24680 a year later, they did not charge anything (no service charge) if you borrowed 24680 and paid back 24681 then you paid them back the original plus an extra $1. the extra $1 is the service charge. what about here?

OpenStudy (anonymous):

a. $396 b. $8,488 c. $8,884 d. $8,292

OpenStudy (anonymous):

these are the answer choices

OpenStudy (phi):

the finance charge is the amount over the original $24,680

OpenStudy (anonymous):

thank u cutie !!

OpenStudy (anonymous):

$33168-($24680+$396)=$8884

OpenStudy (anonymous):

You are considering taking out one of two loans. Loan R has a principal of $17,550, an interest rate of 5.32% (compounded monthly), and a duration of seven years. Loan S has a principal of $15,925, an interest rate of 6.07% (compounded monthly), and a duration of nine years. Assuming that you pay back each in monthly intervals, which loan will have a greater lifetime total, and how much greater will it be? Round all dollar values to the nearest cent.

OpenStudy (anonymous):

oh wait.. if phi said total out is $33564 then subtract ($24680+$396=$25076) = $8488

OpenStudy (anonymous):

ohh lol thanks

OpenStudy (anonymous):

could u help me with the second question

OpenStudy (anonymous):

oh and 8488 was wrong

OpenStudy (anonymous):

8884 is wrong it seems

OpenStudy (anonymous):

8488 is what i get

OpenStudy (anonymous):

8884 might be correct

OpenStudy (anonymous):

i'm not sure phi seems to confused me

OpenStudy (anonymous):

33564-24680=8884

OpenStudy (anonymous):

yeah me to lol .. i will try 8884 and let u kno

OpenStudy (anonymous):

what about the second one ?? any idea ?

OpenStudy (anonymous):

Loan R: $17,550(1+0.0532/12)^12*7=$25,447.84071 Loan S: $15,925(1+0.0607/12)^12*9=$27,462.27 Loan S-Loan R: $27,462.27-$25,447.84071=$2014.43 (amount greater) So loan S would have a greater amount lifetime total, and be greater by $2014.43.

OpenStudy (anonymous):

yes n thank u !! and th 8884 was the right answer for that other question

OpenStudy (anonymous):

there's answers?

OpenStudy (anonymous):

no it was a test

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

yeah lol thanks .. im doing personal loans

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

could u help me ?